OpenStudy (anonymous):

HELP PLEASE ! (: I'll give medals . Find the y-intercepts of the following quadratic function: f(x)= 1/3(x+6)^2-10

OpenStudy (anonymous):

y intercept is when x=0 so set x=0 f(x)=1/3(0+6)^2 - 10 =(1/3)(36)-10 =12-10 =2 at y=2 is ur y intercept (pt 0,2)